OpenStudy (anonymous):

16.61+10y=x sovle so its x= one number

OpenStudy (jessereyna):

Hey

OpenStudy (anonymous):

hia

OpenStudy (jessereyna):

Set the equation equal to y

OpenStudy (jessereyna):

Then set y equal to zero

OpenStudy (anonymous):

i dont know how

OpenStudy (landon34):

Simplifying 16.61 + 10y = x Solving 16.61 + 10y = x Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Add '-16.61' to each side of the equation. 16.61 + -16.61 + 10y = -16.61 + x Combine like terms: 16.61 + -16.61 = 0.00 0.00 + 10y = -16.61 + x 10y = -16.61 + x Divide each side by '10'. y = -1.661 + 0.1x Simplifying y = -1.661 + 0.1x